w=Cr to the power of -2
c is the constant and r is the distance that the object is from the center of the earth.
object weighs 100 pounds. find value of C that makes the equation true

Answer by stanbon(75887)

w=Cr^(-2)
c is the constant and r is the distance that the object is from the center of the earth.
object weighs 100 pounds. find value of C that makes the equation true
---
C = w/r^(-2)
C = 100*r^2
You have to know the radius of the earth to find "C".
Cheers,
Stan H.
=============